OpenStudy (anonymous):

how do you put y-4=2(x-2) and y-8=2(x-4) into slope-intercept form.

OpenStudy (lucaz):

intercept form is y=mx+b

OpenStudy (anonymous):

I know that.. but I have to put those to equations into intercept form... how do I do that?

OpenStudy (lucaz):

just simplify the equations

OpenStudy (lucaz):

hm, wait..

OpenStudy (lucaz):

I think it's y=2x+0 both of them

OpenStudy (texaschic101):

y - 4 = 2(x - 2) -- distribute through the parenthesis y - 4 = 2x - 4 -- add 4 to both sides y = 2x + 0 y - 8 = 2(x - 4) -- distribute through the parenthesis y - 8 = 2x - 8 -- add 8 to both sides y = 2x + 0 I agree with lucaz :)

OpenStudy (anonymous):

simply solve for y

OpenStudy (anonymous):

since it is defined to be y=mx+b all you have to do is solve for y like @texaschic101 did

OpenStudy (texaschic101):

:)